Group Selection Page HELP Controls Group ITDM Vacuum WP78 SPB SQS XO ToBeMigrated2XWiki ToBeMigrated2GitlabPages ToBeMigratedByGroup Migrated2Zulip Archived2PDF not2MigrateOrArchive
SCS SXP Migrated2Zulip Photon Commissioning Team XRO DOC HED Migrated2Zulip WP74 MID PSPO_XO XO_Operation SQS Migrated2Zulip FXE Migrated2Zulip DET Sample Environment SPB Migrated2Zulip
OPERATION-2025 OPERATION-2024 OPERATION-2023 OPERATION-2022 OPERATION-2021 OPERATION-2020 OPERATION-2019 SASE1-2018 SASE1-2017 Experiment data
  OPERATION-2025 logbook, Contact: F.Wolff-Fabris, A.Galler, H.Sinn  Not logged in ELOG logo
Message ID: 58     Entry time: 30 Apr 2025, 22:46
Author: Romain Letrun 
Group: SPB 
Subject: Shift summary 

X-ray delivery

  • 9.3 keV, 1800 uJ pulse energy
  • 1 - 202 pulses at 564 kHz

  • Poor beam stability overall. Lots of intensity fluctuations and rather long time to recover after change of number of pulses in any beamline.

Optical laser delivery

  • 800 nm, 113 pulses, 564 kHz stable delivery

  • Illumination laser aligned

Achievements/Observations

  • Aligned instrument

  • Prepared illumination laser
  • Shot PCS sample for a few minutes

Issues

  • PCS sample clogged (at filter not nozzle) after a few crystal shots

  • A few clogs with nozzle with lyzo, when X-ray were on, even far from the nozzle somehow the splash from the explosion grew on the tip

ELOG V3.1.4-7c3fd00